Grade 12General Physics

A planet of mass m( much less than that of sun) is in an elliptical orbit about the sun with an orbital period T. If A be the area of the orbit, what is it's angular momentum?

divyan khare
12 Years agoGrade 12

1 Answer

Nirmal Singh.
12 Years ago

we know the formula of angular momentum = m*v*r----------------(i)-

we also know that v = 2*pi*r/T

and by kepler's law we know know T2 = (4*pi^2/GM)*r^3

Consider M is the mass of Sun r = (GMT^2/4 pi^2)^1/3

now we will put all value in (i)

and we will get the answer

Regards,

Nirmal Singh

Askiitians faculty